OpenStudy (anonymous):

Solve for x, and write the exact answer using base-10 logarithms. 2^(9x)=16^(-x+9)

OpenStudy (anonymous):

2^(9x)=(2^4)^(-x+9) 2^(9x)=2^(-4x+36) 9x=-4x+36 13x=36 x=36/13

OpenStudy (anonymous):

what does this have to do with base ten logarithms? or any logarithms?

OpenStudy (anonymous):

i dont know, thats why i'm asking!